# FIDLAR: Forecast-Informed Deep Learning Architecture for Flood Mitigation

### v1

@@ C.2 Flood mitigation at all locations
In Table 3, we provided the comparison of the performance of different models for the Flood Manager at time point $t+1$, but at one location (S1). We provide the comparison results of different models for the Flood Manager on the test set (at time t+1 for all locations) in Table 9.
@@
GA-based Genetic $*{}^{*}$ (GA) - - - -
Genetic ${}^{{\dagger}}$ (GA) 343 67.50 1,817 416
@@
${}^{{\ddagger}}$ (Ours) 114 12.09 617 98.42
Table 9: Comparison of different methods for the Flood Manager on the test set (at time t+1 for all stations). The $*{}^{*}$ denotes that the GA method was used for Flood Manager with a physics-based (HEC-RAS) as an Evaluator. -- denotes too long to get results. The ${}^{{\dagger}}$ denotes that the GA method was used for Flood Manager with the DL-based GTN as the Evaluator. All other rows are DL-based flood managers with a DL-based GTN as the Evaluator. ${}^{{\ddagger}}$ denotes the best.
@@ C.3 Flood mitigation with one event
In Fig. 7, we visualize the results of different models for the Flood Manager on a short event spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest. Below are the corresponding experiment results.
@@
GA-Based Genetic $*{}^{*}$ (GA) 4 0.351 0 0
Genetic ${}^{{\dagger}}$ (GA) 6 0.764 0 0
@@
${}^{{\ddagger}}$ (Ours) 1 0.022 0 0
Table 10: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event S1).
@@ Appendix D Visualization
Fig. 11 presents the results for all locations compared to Fig. 7 for one location of interest. Table 11 shows the corresponding performance measures.
@@
GA-Based Genetic $*{}^{*}$ (GA) 16 1.39 0 0
Genetic ${}^{{\dagger}}$ (GA) 23 3.75 0 0
@@
${}^{{\ddagger}}$ (Ours) 5 0.22 0 0
Table 11: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event at all locations).

### v2

@@ C.2 Flood mitigation at all locations
In Table 3, we provided the comparison of the performance of different models for the Flood Manager at time point $t+1$, but at one location (S1). We provide the comparison results of different models for the Flood Manager on the test set (at time t+1 for all locations) in Table 9. All DL-based methods show better results than “rule-based” and GA-based approaches for flood mitigation. Our GTN model provides the best performance in flood mitigation and comparable performance in avoiding water wastage.
Table 9: Comparison of different methods for the Flood Manager on the test set (at time t+1 for all stations). The denotes that the GA method was used for Flood Manager with a physics-based (HEC-RAS) as an Evaluator. -- denotes too long to get results. The denotes that the GA method was used for Flood Manager with the DL-based GTN as the Evaluator. All other rows are DL-based flood managers with a DL-based GTN as the Evaluator. We mark the best results in bold.
@@
GA-based Genetic Algorithm - - - -
Genetic Algorithm 343 67.50 1,817 416
@@
GTN (Ours) 114 12.09 617 98.42
@@ C.3 Flood mitigation with one event
In Figure 7, we visualize the results of different models for the Flood Manager on a short event spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest. Below are the corresponding experiment results. Our GTN shows the best performance in flood mitigation.
Table 10: Comparison of different Managers for flood Mitigation (at time t+1 for one event S1). The best results are in bold.
@@
GA-Based Genetic Algorithm (GA) 4 0.351 0 0
Genetic Algorithm (GA) 6 0.764 0 0
@@
GTN (Ours) 1 0.022 0 0
@@ Appendix D Visualization
Figure 10 presents the results for all locations compared to Figure 7 for one location of interest. Table 11 shows the corresponding performance measures. Our GTN shows the best performance in flood mitigation.
@@
Table 11: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event at all locations). We mark the best results in bold.
@@
GA-Based Genetic Algorithm (GA) 16 1.39 0 0
Genetic Algorithm (GA) 23 3.75 0 0
@@
GTN (Ours) 5 0.22 0 0